OpenStudy (anonymous):

When 36 is subtracted of a number, the result is five times the number. What is the positive solution? 1) 9 2) 6 3) 3 4) 4

OpenStudy (anonymous):

from the square of a number*

OpenStudy (anonymous):

When 36 is subtracted from a square of a number, the result is five times the number. What is the positive solution?

OpenStudy (ash2326):

9 x^2-36=5x x^2-5x-36=0 x^2-9x+4x-36=0 x(x-9)+4(x-9)=0 (x-9)(x+4)=0 x=9 or -4 the positive solution is 9

OpenStudy (king):

the number is 4 or -9

OpenStudy (king):

so positive solution is 9 sry its -4

OpenStudy (anonymous):

its ok lol